# Rest energy effects for non-relativistic Fermi fields in external potentials

### v3

@@
Abstract
@@
The generalized Lévy-Leblond equation (GLL) is used to study the bound state problem for a non-relativistic Fermi field in external potentials. A condition when the rest energy term can be removed from the GLL equation is determined. For external potentials which do not obey the condition, the bound state spectrim is affected non-trivially by the rest energy of the Fermi field. This is demonstated for a spherical well type potential of finite depth. The existence of bound states in the well, their number and energies all depend on the value of the rest energy.
I Introduction
@@
Non-relativistic particles (fields), defined through the theory of representations of the Galilei group, are known to acquire spin and magnetic moment levy , the properties which were originally believed to be purely relativistic. The dynamics of the spin $1/2$ particles is governed by the Lévy-Leblond (LL) equation, the non-relativistic version of the Dirac equation. In levy , it was derived in the Dirac way by linearization of the Schrödinger equation.
@@
Another way of deriving the LL equation is by making use of a $(4+1)$-dimensional covariant formulation of Galilean covariance ap1999 based on an extended space-time approach takahashi ,omote . The idea of such approach in the non-relativistic framework appeared first in eisen and then rediscovered in duval1985 . It was used to study the symmetries of the LL equation in the non-relativistic Chern-Simons electrodynamics duval1996 . A formulation of Galilean covariance in one higher dimension allows us to incorporate into the theory the notion of rest (internal) energy: one of the Casimir invariants of the Galilei group in $(4+1)$-dimensions includes the rest energy operator. Reduction to $(3+1)$-dimensions by eliminating the additional, fifth coordinate results in the generalized LL equation with a rest energy contribution sar2011 . A generalized version of the LL equation was also discussed in hueg2013 .
@@
It was shown in sour that the rest energy is an additive constant related to the particular structure of the Galilei group and that the relativistic energy does not contain such constant. Any additive energy constant including the rest energy can be considered as a reference point for the energy of non-relativistic systems, and the only effect of taking this constant into account is a shift of all energy levels by its value. This is equivalent to a change in a phase of wave function. In this paper, we study the GLL equation in external potentials, and we aim to show that for some external potentials the additivity of the rest energy is violated. The rest energy term cannot be removed by a shift of energy levels or by redefining the Fermi field components. This results in a non-trivial impact of the rest energy of non-relativistic Fermi particles on their dynamics and, in particular, on the bound state spectrum.
@@
Our paper is organized as follows. In Sect.2, we consider a free non-relativistic Fermi field. We define a transformation of the Fermi field which removes the rest energy term from the GLL equation. The transformation is a combination of phase change and mixing the field componenets. This demonstrates the equivalence of the GLL and LL equations for free non-relativistic Fermi fields.
@@
We establish a two-step transition from the Dirac equation to the GLL equation. Taking the non-relativistic limit of the Dirac equation yields the LL equation levy , and an additional transformation is required to get its generalized version.
@@
In Sect.3, we consider a non-relativistic Fermi field in external potentials. We find that for some types of potentials the rest energy cannot be removed from the GLL equation. The condition when the removal can be done is determined. The component form of the GLL equation with an external potential is given.
@@
As an example of external potential for which the condition does not hold, we study in detail the case of a potential of spherical well type of finite depth. We derive a Schrödinger type equation for the independent component of the Fermi field with an effective potential exhibiting the rest energy dependance. We determine the spectrum of the bound S-states, and we find the minimum value of the rest energy below which no bound states exist.
@@
II GLL equation for a free non-relativistic Fermi field
@@
By re-defining the wave function ${\psi}_{+}({\bf r},t)$ as
@@
The GLL equation for a $4$-component non-relativistic Fermi field ${\psi}({\bf r},t)$ reads sar2011
@@
$I_{+}\equiv I-\frac{m{c}}{k}{\gamma}^{5},$
@@
fulfil the algebra
$\gamma^{\mu}\gamma^{\nu}+\gamma^{\nu}\gamma^{\mu}=2g^{\mu\nu},$
@@
$g^{\mu\nu}=\left(\begin{array}[]{ccc}-{\mathbf{1}}_{3\times 3}&0&0\\ 0&0&1\\ 0&1&0\end{array}\right).$
@@
${\psi}({\bf r},t)\to{\eta}({\bf r},t)\equiv\left(I-\frac{{\mu}}{\sqrt{2}}{\gamma}^{4}\right)e^{\frac{\rm i}{\hbar}E_{\rm rest}t}{\psi}({\bf r},t),$
@@
$\Big{[}{\rm i}{\hbar}\gamma^{4}\partial_{4}+{\rm i}{\hbar}\left(I-\frac{{\mu}}{\sqrt{2}}{\gamma}^{4}\right){\gamma}^{a}{\partial}_{a}$
$-kI_{+}\left(I+\frac{{\mu}}{\sqrt{2}}{\gamma}^{4}\right)+\frac{1}{c}E_{\rm rest}{\gamma}^{4}\Big{]}{\eta}({\bf r},t)=0.$
@@
$\Big{[}{\rm i}{\hbar}\gamma^{\bar{\mu}}\partial_{\bar{\mu}}+m{c}{\gamma}^{5}+\left({\mu}m{c}\sqrt{2}-k\right)I$
$+\frac{1}{2m{c}}{\left({\mu}m{c}\sqrt{2}-k\right)}^{2}{\gamma}^{4}\Big{]}{\eta}({\bf r},t)=0.$ (14)
@@
${\eta}({\bf r},t)=\left(I-\sqrt{\frac{m_{0}}{2m}}{\gamma}^{4}\right)e^{\frac{\rm i}{\hbar}E_{\rm rest}t}{\psi}({\bf r},t),$ (16)
@@
${\eta}({\bf r},t)=\left(\begin{array}[]{c}{\eta}_{1}({\bf r},t)\\ {\eta}_{2}({\bf r},t)\end{array}\right),$
@@
The GLL equation for a free non-relativistic Fermi field can be therefore reduced to the Schrödinger equation for its independent component. In this sense, the GLL equation and the Schrödinger equation are equivalent to each other. However, starting with the GLL equation allows us to study both the spin and rest energy effects.
The GLL equation can be connected with the Dirac equation as well. The rest energy of a free relativistic Fermi field is incorporated nontrivially in its total or relativistic energy. Its rest mass $m_{0}$ is at the same time a measure of its inertia, i.e. of the tendency of the field to resist changes in velocity. Thus, $m_{0}$ is its inertial mass as well jam .
For non-relativistic Fermi fields, the rest energy contribution decouples from the kinetic energy and becomes an additive factor. In this case, the rest energy is independent from inertia, and the rest mass $m_{0}$ is not in general the same as the inertial mass $m$ jack .
@@
As a result, the non-relativistic limit of the Dirac equation is the LL equation without any rest energy contribution levy . It is sometimes stated that the rest energy is a purely relativistic effect and has no non-relativistic limit tha .
@@
III GLL equation with external potentials
@@
$\tilde{V}\equiv V+\frac{1}{\sqrt{2}}{\mu}\left\{{\gamma}^{4},V\right\}+\frac{1}{2}{\mu}^{2}{\gamma}^{4}\left\{{\gamma}^{4},V\right\}$ (24)
@@
${\rm i}{\hbar}{\gamma}^{4}{\partial}_{4}\to\frac{1}{c}{\gamma}^{4}\left({\rm i}{\hbar}\frac{\partial}{{\partial}t}-{\varphi}\right),$
indicating that ${\varphi({\bf r},t)}$ in Eq.(26) can be interpreted as potential energy (or potential function) of the system.
Another example is a potential produced by an electromagnetic field $A_{\bar{\mu}}$. It can be introduced by replacing ${\partial}_{\bar{\mu}}$ with $D_{\bar{\mu}}\equiv{\partial}_{\bar{\mu}}-\frac{\rm i}{\hbar}\frac{e}{c}A_{\bar{\mu}}$, where $e$ is the charge of the Fermi field. This yields
@@
Since the matrices ${\gamma}^{\bar{\mu}}$ anti-commute with ${\gamma}^{4}$, the condition (25) is satisfied and the rest energy $E_{\rm rest}$ does not show itself in the solutions of Eq.(23). The GLL equation with the potential given by Eq.(27) is the linearized version of the Pauli equation gre ,wilk .
@@
$\tilde{V}=\frac{1}{c}\left(I+{\mu}\sqrt{2}{\gamma}^{4}\right){\varphi}$
@@
$\bar{\varphi}({\xi})=\left\{\begin{array}[]{cc}-\bar{\varphi}_{0}&{\rm for}\hskip 14.22636pt{\xi}<{\xi}_{0}\\ 0&{\rm for}\hskip 14.22636pt{\xi}>{\xi}_{0},\end{array}\right.$ (30)
@@
$V_{\rm eff}=\left\{\begin{array}[]{cc}-V_{0}&{\rm for}\qquad{\xi}<{\xi}_{0}\\ 0&{\rm for}\qquad{\xi}>{\xi}_{0},\end{array}\right.$ (33)
@@
$V_{0}={\mu}\sqrt{2}\bar{\varphi}_{0}-\frac{1}{2}{\bar{\varphi}_{0}}^{2},$
@@
In what follows, we assume that ${\mu}>{\mu}_{0}$, so that we can search for bound states. The spin terms do not appear in the effective Hamiltonian density. The set of operators commuting with ${\cal H}$ is ${\bf L}^{2}$,${\rm L}_{3}$, so the solutions of the eigenvalue equation (31) have the form
@@
${\rm L}_{3}{\eta}_{+,lM}$ $=$ $M{\eta}_{+,lM}.$
Substituting the ansatz (36) into the eigenvalue equation (31) and using the expression for ${\bf p_{\xi}^{2}}$ in spherical polar coordinates
${\bf p_{\xi}^{2}}=-\frac{{\partial}^{2}}{{\partial}{\xi}^{2}}-\frac{2}{\xi}\frac{\partial}{{\partial}{\xi}}+\frac{1}{{\xi}^{2}}{\bf L}_{\xi}^{2},$
@@
We are interested in the bound state solutions for which $|{\cal E}|<V_{0}$. The solutions are flugge ,capri
$R_{l}({\xi})=\left\{\begin{array}[]{cc}N_{-}j_{l}(a{\xi})&{\rm for}\qquad{\xi}<{\xi}_{0}\\ N_{+}h_{l}^{(1)}({\rm i}b{\xi})&{\rm for}\qquad{\xi}>{\xi}_{0},\end{array}\right.$ (40)
@@
${\rho}\equiv\sqrt{1-\frac{|{\cal E}|}{V_{0}}},\qquad\bar{\xi}_{0}\equiv\sqrt{2V_{0}}{\xi}_{0}.$
@@
${\mu}\geq{\mu}_{\rm min}\equiv{\mu}_{0}+\frac{{\pi}^{2}}{8\sqrt{2}\bar{\varphi}_{0}{\xi}_{0}^{2}},$
can Eq.(42) have a solution. Herein, ${\mu}_{\rm min}$ is the minimum non-zero value of ${\mu}$ to achieve the bound S-states, and
$E_{0,{\rm min}}\equiv\frac{1}{8}{\left[\bar{\varphi}_{0}+\frac{{\pi}^{2}}{4\bar{\varphi}_{0}{\xi}_{0}^{2}}\right]}^{2}mc^{2}$
is the corresponding minimum value of rest energy.
IV Discussion

### v4

@@
Rest energy effects for non-relativistic Fermi fields in external potentials
Fuad M. Saradzhev 111fuads@athabascau.ca Centre for Science, Athabasca University, Athabasca, Alberta, Canada
@@
The generalized Lévy-Leblond equation (GLL) is used to study the bound state problem for a non-relativistic Fermi field in external potentials. A condition when the rest energy term can be removed from the GLL equation is determined. For external potentials which do not obey the condition, the bound state spectrum is affected non-trivially by the rest energy of the Fermi field. This is demonstrated for a spherical finite-depth well potential. The existence, number, and energies of bound states depend on the value of the rest energy.
Keywords: Lévy-Leblond equation, Fermi fields, Bound states, Rest energy
1 Introduction
@@
Non-relativistic particles (fields), defined through the theory of representations of the Galilei group, are known to acquire spin and magnetic moment [1]-[3], the properties which were originally believed to be purely relativistic. The dynamics of the spin $1/2$ particles is governed by the Lévy-Leblond (LL) equation, the non-relativistic version of the Dirac equation. In [1], it was derived in the Dirac way by linearization of the Schrödinger equation.
@@
Another way of deriving the LL equation is by making use of a $(4+1)$-dimensional covariant formulation of Galilean covariance [4]-[11] based on an extended space-time approach [12]-[14]. The idea of such an approach in the non-relativistic framework appeared first in [15] and then rediscovered in [16]. It was used to study the symmetries of the LL equation in the non-relativistic Chern-Simons electrodynamics [17]. A formulation of Galilean covariance in one higher dimension allows us to incorporate into the theory the notion of rest (internal) energy: one of the Casimir invariants of the Galilei group in $(4+1)$-dimensions includes the rest energy operator. Reduction to $(3+1)$ dimensions by eliminating the additional, fifth coordinate results in the generalized LL equation with a rest energy contribution [18]. A generalized version of the LL equation was also discussed in [19].
@@
It was shown in [20] that the rest energy is an additive constant related to the particular structure of the Galilei group and that the relativistic energy does not contain such constant. Any additive energy constant including the rest energy can be considered as a reference point for the energy of non-relativistic systems, and the only effect of taking this constant into account is a shift of all energy levels by its value. This is equivalent to a phase change in the wave function. In this paper, we study the GLL equation in external potentials, and we aim to show that for some external potentials the assumption of rest energy additivity is violated. The rest energy term cannot be removed by a shift of energy levels or by redefining the Fermi field components.
The physical consequences of rest energy in non-relativistic physics have not been studied before. Our motivation is to reveal the dynamical effects of rest energy, in particular, its non-trivial impact on the bound state spectrum of non-relativistic Fermi particles.
@@
Our paper is organized as follows. In Sect.2, we consider a free non-relativistic Fermi field. We define a transformation for the Fermi field which removes the rest energy term from the GLL equation. The transformation is a combination of phase change and mixing the field components. This demonstrates the equivalence of the GLL and LL equations for free non-relativistic Fermi fields.
@@
We establish a two-step transition from the Dirac equation to the GLL equation. Taking the non-relativistic limit of the Dirac equation yields the LL equation [1], and an additional transformation is required to get its generalized version.
@@
In Sect.3, we consider a non-relativistic Fermi field in external potentials. We find that for some types of potentials the rest energy cannot be removed from the GLL equation. We determine the condition under which the rest energy can be removed. The component form of the GLL equation with an external potential is given.
@@
As an example of external potential for which the condition does not hold, we study in detail the case of a potential of spherical well type of finite depth. We derive a Schrödinger type equation for the independent component of the Fermi field with an effective potential exhibiting the rest energy dependence. We determine the spectrum of the bound S-states, and we find the minimum value of the rest energy below which no bound states exist.
@@
2 GLL equation for a free non-relativistic Fermi field
@@
By re-defining the wave function ${\psi}({\bf r},t)$ as
@@
The GLL equation for a $4$-component non-relativistic Fermi field ${\psi}({\bf r},t)$ reads [18]
@@
$I_{+}\equiv I-\frac{m{c}}{k}{\gamma}^{5},$ (6)
@@
satisfy the algebra
$\gamma^{\mu}\gamma^{\nu}+\gamma^{\nu}\gamma^{\mu}=2g^{\mu\nu},$ (15)
@@
$g^{\mu\nu}=\left(\begin{array}[]{ccc}-{\mathbf{1}}_{3\times 3}&0&0\\ 0&0&1\\ 0&1&0\end{array}\right).$ (16)
@@
${\psi}({\bf r},t)\to{\eta}({\bf r},t)\equiv\left(I-\frac{{\mu}}{\sqrt{2}}{\gamma}^{4}\right)e^{\frac{\rm i}{\hbar}E_{\rm rest}t}{\psi}({\bf r},t),$ (17)
@@
$\Big{[}{\rm i}{\hbar}\gamma^{4}\partial_{4}+{\rm i}{\hbar}\left(I-\frac{{\mu}}{\sqrt{2}}{\gamma}^{4}\right){\gamma}^{a}{\partial}_{a}-kI_{+}\left(I+\frac{{\mu}}{\sqrt{2}}{\gamma}^{4}\right)+\frac{1}{c}E_{\rm rest}{\gamma}^{4}\Big{]}{\eta}({\bf r},t)=0.$ (18)
@@
$\Big{[}{\rm i}{\hbar}\gamma^{\bar{\mu}}\partial_{\bar{\mu}}+m{c}{\gamma}^{5}+\left({\mu}m{c}\sqrt{2}-k\right)I+\frac{1}{2m{c}}{\left({\mu}m{c}\sqrt{2}-k\right)}^{2}{\gamma}^{4}\Big{]}{\eta}({\bf r},t)=0.$ (19)
@@
${\eta}({\bf r},t)=\left(I-\sqrt{\frac{m_{0}}{2m}}{\gamma}^{4}\right)e^{\frac{\rm i}{\hbar}E_{\rm rest}t}{\psi}({\bf r},t)$ (21)
@@
${\eta}({\bf r},t)=\left(\begin{array}[]{c}{\eta}_{1}({\bf r},t)\\ {\eta}_{2}({\bf r},t)\end{array}\right),$ (23)
@@
The GLL equation for a free non-relativistic Fermi field can be therefore reduced to the $2$ x $2$ matrix Schrödinger equation for its independent component. In this sense, the GLL equation and the Schrödinger equation are equivalent to each other. However, starting with the GLL equation allows us to study both the spin and rest energy effects.
The GLL equation can be connected with the Dirac equation as well. The rest energy of a free relativistic Fermi field is incorporated nontrivially in its total or relativistic energy. Its rest mass $m_{0}$ is at the same time a measure of its inertia, i.e. of the tendency of the field to resist changes in velocity. Thus, $m_{0}$ is its inertial mass as well [22].
For non-relativistic Fermi fields, the rest energy contribution decouples from the kinetic energy and becomes an additive factor. In this case, the rest energy is independent from inertia, while the rest mass $m_{0}$ is not in general the same as the inertial mass $m$ [23],[24] and can be interpreted as a parameter related to intrinsic structure of Fermi particles.
@@
As a result, the non-relativistic limit of the Dirac equation is the LL equation without any rest energy contribution [1]. It is sometimes stated that the rest energy is a purely relativistic effect and has no non-relativistic limit [25].
@@
3 GLL equation with external potentials
@@
$\tilde{V}\equiv V+\frac{1}{\sqrt{2}}{\mu}\left\{{\gamma}^{4},V\right\}+\frac{1}{2}{\mu}^{2}{\gamma}^{4}\left\{{\gamma}^{4},V\right\},$ (31)
@@
${\rm i}{\hbar}{\gamma}^{4}{\partial}_{4}\to\frac{1}{c}{\gamma}^{4}\left({\rm i}{\hbar}\frac{\partial}{{\partial}t}-{\varphi}\right),$ (34)
indicating that ${\varphi({\bf r},t)}$ in Eq.(33) can be interpreted as potential energy (or potential function) of the system. For the hydrogen atom, ${\varphi}$ is the potential energy of the Coulomb interaction with no rest energy contribution [26], [27].
The condition (32) also holds for the potential produced by an electromagnetic field $A_{\bar{\mu}}$. It can be introduced by replacing ${\partial}_{\bar{\mu}}$ with $D_{\bar{\mu}}\equiv{\partial}_{\bar{\mu}}-\frac{\rm i}{\hbar}\frac{e}{c}A_{\bar{\mu}}$, where $e$ is the charge of the Fermi field. This yields
@@
The matrices ${\gamma}^{\bar{\mu}}$ anti-commute with ${\gamma}^{4}$, and the rest energy $E_{\rm rest}$ does not show itself in the solutions of Eq.(30). The GLL equation with the potential given by Eq.(35) is the linearized version of the Pauli equation [2],[28].
@@
$\tilde{V}=\frac{1}{c}\left(I+{\mu}\sqrt{2}{\gamma}^{4}\right){\varphi}$ (37)
@@
$\bar{\varphi}({\xi})=\left\{\begin{array}[]{cc}-\bar{\varphi}_{0}&{\rm for}\hskip 14.22636pt{\xi}<{\xi}_{0},\\ 0&{\rm for}\hskip 14.22636pt{\xi}>{\xi}_{0},\end{array}\right.$ (41)
@@
$V_{\rm eff}=\left\{\begin{array}[]{cc}-V_{0}&{\rm for}\qquad{\xi}<{\xi}_{0},\\ 0&{\rm for}\qquad{\xi}>{\xi}_{0},\end{array}\right.$ (44)
@@
$V_{0}={\mu}\sqrt{2}\bar{\varphi}_{0}-\frac{1}{2}{\bar{\varphi}_{0}}^{2},$ (45)
@@
In what follows, we assume that ${\mu}>{\mu}_{0}$, so that we can search for bound states. The spin terms do not appear in the effective Hamiltonian density. The set of operators commuting with ${\cal H}$ is ${\bf L}^{2}$,${\rm L}_{3}$, so the solutions of the Eq.(42) have the form
@@
${\rm L}_{3}{\eta}_{+,lM}$ $=$ $M{\eta}_{+,lM}.$ (49)
Substituting the ansatz (48) into Eq.(42) and using the expression for ${\bf p_{\xi}^{2}}$ in spherical polar coordinates
${\bf p_{\xi}^{2}}=-\frac{{\partial}^{2}}{{\partial}{\xi}^{2}}-\frac{2}{\xi}\frac{\partial}{{\partial}{\xi}}+\frac{1}{{\xi}^{2}}{\bf L}_{\xi}^{2},$ (50)
@@
We are interested in the bound state solutions for which $|{\cal E}|<V_{0}$. The solutions are [29],[30]
$R_{l}({\xi})=\left\{\begin{array}[]{cc}N_{-}j_{l}(a{\xi})&{\rm for}\qquad{\xi}<{\xi}_{0},\\ N_{+}h_{l}^{(1)}({\rm i}b{\xi})&{\rm for}\qquad{\xi}>{\xi}_{0},\end{array}\right.$ (54)
@@
${\rho}\equiv\sqrt{1-\frac{|{\cal E}|}{V_{0}}},\qquad\bar{\xi}_{0}\equiv\sqrt{2V_{0}}{\xi}_{0}.$ (57)
@@
Figure 3: ${\mu}-\bar{\varphi}_{0}$ phase diagram for the effective potential $V_{\rm eff}$ with ${\xi}_{0}=2.5$. There are no bound states in phase I, where $V_{\rm eff}$ is a potential step, and in phase II, where $V_{\rm eff}$ is a spherical well potential but the Fermi particle does not have enough rest energy to be bounded. The phase with bound ${\rm S}$-states is phase III.
@@
${\mu}\geq{\mu}_{\rm min}\equiv{\mu}_{0}+\frac{{\pi}^{2}}{8\sqrt{2}\bar{\varphi}_{0}{\xi}_{0}^{2}}$ (60)
can Eq.(56) have a solution. Herein, ${\mu}_{\rm min}$ is the minimum non-zero value of ${\mu}$ to achieve the bound S-states (see Figure 3), and
$E_{{\rm rest},{\rm min}}\equiv\frac{1}{8}{\left[\bar{\varphi}_{0}+\frac{{\pi}^{2}}{4\bar{\varphi}_{0}{\xi}_{0}^{2}}\right]}^{2}mc^{2}$ (61)
is the corresponding minimum value of rest energy. For an electron, the estimates of $E_{{\rm rest},{\rm min}}$ are $4.14$ MeV (for $\bar{\varphi}_{0}=8$, ${\xi}_{0}=2.5$) and $0.31$ MeV (for $\bar{\varphi}_{0}=2$, ${\xi}_{0}=2.5$).
4 Discussion